a rectangle is 6 inches longer than it is wide the area of the reran gel is 112 square inches find the length and width of the rectangle

Answer by Boreal(15235)

L*w=Area=112 in^2
w
L=w +6
w(w+6)=112
w^2+6w-112=0 This is distributing the w and subtracting 112 from both sides. I now have a quadratic equation.
Factor it
4/23 no 8/14 yes
(w+14)(w-8)=0 ; This works if either w+14 or w-6=0
w=-14, w=8 ; only the second makes sense
length=8+6=14
14*8=112 (in*in=in^2).
The length is 14 and the width is 8